OpenStudy (anonymous):

How to simplify this expression: (2x/x^2-6x+9) - (1/x+1) - (8/x^2-2x-3)

OpenStudy (anonymous):

\[\frac{ 2x }{ x^2-6x+9 } - \frac{ 1 }{ x+1 } - \frac{ 8 }{ x^2-2x-3 }\]

OpenStudy (blurbendy):

2x / x^2 = 2/x (2/x - 6x + 9) -(1/x +1) - (8/x^2 -2x -3) Next distribute the negative signs in front of (1/x + 1) and -(8/x^2 -2x -3)

zepdrix (zepdrix):

The denominators of the first and last term can be factored, have you tried doing that yet? :D

OpenStudy (anonymous):

I was able to solve it. Thanks!